A Force F= (6xi+ 5yj), where F is in newtons and x and y are in meters, acts on an object as the object moves in the x direction from the origin to x = 4.69 m.

Find the work  W =  integral (F
Answer #1

65.9883 J
Answer #2

Work is found with a vector calculus line integral
W = ?F
Answer #3

work will be done because of x component of the force.


W = integral 6x dx = 12 *x^2 / 2 = 6 x^2 = 6 X 4.69 X 4.69 = 131.98 J
Answer #4

vector r = 4.69i

now work = Integral of (6xi+ 5yj).4.69i dx = Integral of 28.14 x.dx from 0 to 4.69

W = (28.14 x^2)/2 = 14.07x^2

putting value of x = 4.69

W = 309.48 J
Answer #5

Work is found with a vector calculus line integral
W = ?F
Answer #6

Work is found with a vector calculus line integral
W = ?F